# Fast constrained sampling in pre-trained diffusion models

### v2

@@ C.3 Inference visualizations
C.4 Effect of convergence speed on final images
We ask the question of how does convergence speed affect the quality of the generated images? Previous works found that applying a high weight on the constraint led to unwanted artifacts in the generated images (Chung et al., 2023). We hypothesize that, apart from artifacts in the gradient, ’over-optimizing’ for the condition at a given timestep can affect the generation quality. In practice, if we push the initial ${\bm{x}}_{t}$ too far from the inputs the denoiser network is expecting, either with a high weight on the gradient update or by performing too many updates, we should be seeing non-realistic images in the output.
@@ C.4 Effect of convergence speed on final images
Figure 14: Convergence for the inpainting task when using 5 optimization steps $K$.
Figure 15: Convergence for the inpainting task when using 20 optimization steps $K$.
C.5 More Qualitative Results
@@
In Figure 16 we showcase additional results on the box inpainting task. MPGD (He et al., 2024), which does not backpropagate the constraint error through the diffusion model completely fails at inpainting the missing region. We attribute that to the minimal ability to influence pixels that are ’far’ from the constraint at lower noise levels without probing the model weights. The Naive algorithm replaces the known pixels in the estimated final image at every denoising iteration.
@@
In Figure 17 we present additional results on style-guided text-to-image generation for a single prompt. Qualitatively, we see that the style of the images generated with our algorithm better matches the style of the reference image, even when using a different model to define style (OpenCLIP). In Figure 18 we show images generated with different styles and text prompts. Here, we show how increasing the classifier-free guidance weight $w$ (Ho & Salimans, 2022) controls the influence of the text prompt on the final generated image.
@@
In Figure 19 we provide qualitative results of our segmentation mask-guided generation experiment. We see that the baseline of MPGD either over-satisfies the constraint by burning in artifacts ($\rho=1$) or fails to generate images that adhere to the constraint ($\rho=0.5$). Our approach generates the most realistic images while also getting the mask prediction to match to the reference image. Since we are using Stable Diffusion, we used the text prompt ’a headshot photo’ to constrain the set of possible generated images. As expected, the results are not on par with an FFHQ-specific model, but we still find our algorithm able to synthesize good images even in this difficult case.
In Figure 20 we include more samples from the experiment in Section 3.2 in the main paper. We observe that the proposed gradient updates regularly give more consistent outputs that repeat textures across the image. We attribute this to the difference between ${\bm{J}}$ and ${\bm{J}}^{T}$ we discussed in A.2.
Figure 20: Comparison between gradient descent and the proposed gradient. The setup is the same one as in Section 3.2.

### v3

@@ B.5 Inference visualizations
B.6 Effect of convergence speed on final images
We ask the question of how does convergence speed affect the quality of the generated images? Previous works found that applying a high weight on the constraint led to unwanted artifacts in the generated images [5]. We hypothesize that, apart from artifacts in the gradient, ’over-optimizing’ for the condition at a given timestep can affect the generation quality. In practice, if we push the initial ${\bm{x}}_{t}$ too far from the inputs the denoiser network is expecting, either with a high weight on the gradient update or by performing too many updates, we should be seeing non-realistic images in the output.
@@ B.6 Effect of convergence speed on final images
Figure 20: Convergence for the inpainting task when using $K=5$ optimization steps.
Figure 21: Convergence for the inpainting task when using $K=20$ optimization steps.
B.7 More Qualitative Results
@@
In Figure 22 we showcase additional results on the box inpainting task. MPGD [13], which does not backpropagate the constraint error through the diffusion model completely fails at inpainting the missing region. We attribute that to the minimal ability to influence pixels that are ’far’ from the constraint at lower noise levels without probing the model weights. The Naive algorithm replaces the known pixels in the estimated final image at every denoising iteration.
@@
In Figure 23 we present additional results on style-guided text-to-image generation for a single prompt. Qualitatively, we see that the style of the images generated with our algorithm better matches the style of the reference image, even when using a different model to define style (OpenCLIP). In Figure 24 we show images generated with different styles and text prompts. Here, we show how increasing the classifier-free guidance weight $w$ [15] controls the influence of the text prompt on the final generated image.
@@
Appendix C Societal Impact
The work presented in this paper aims to advance the field of machine learning, specifically generative modeling. Solving constrained sampling tasks with a generative prior, can greatly benefit from the better utilization of the image prior. One specific domain is compressed sensing in medical imaging, where generative priors like diffusion models have been used to reconstruct low-dose CT scans and accelerated MRIs. We leave to future work the application of the proposed algorithm to these settings.
However, we acknowledge that there are potential societal consequences of our work that can have a negative impact. The one we highlight is the ability to edit images with the intent to deceive and mislead. While it is true that existing models can already be used to alter images, we understand that our work could offer more precise control over the generation and lead to more convincingly fabricated content.